Electronics > Metrology

HP 5065A Rubidium Standard Repair Log

<< < (17/17)


If you de-flood too long you can overshoot!
If so once you hook the TED back to normal the signal  level will drop slowly till it stabilizes at the new level.
It does sound like you have an offset. Check the zeroing adjustments for A9.
Have you checked the phasing of the 137Hz? Try checking using this procedure.



Corby, thanks for the procedure doc - much appreciated.

In open-loop I set the OCXO fine adjust to just get 273Hz on A7 TP2, although the signal level was only 100mVpp.  I went through the A3 adjustment procedure a few times but can only achieve a max A7 TP2 ac voltage of about 140mVpp, so nowhere near 500mVpp.

With the A8 sync adjustment, I had gone through that before but had not observed the simple half-sine waveform, but rather a higher frequency and noisier signal, but still showing the 'splitting'.  With respect to your notes, the fine freq adjust does vary the level of splitting and that splitting can be effectively nulled (although the phase pot is at full CW), but the 'signal' is now just noise (and A7 TP2 is effectively the same 273Hz amplitude of about 100 to 120mVpp).

With the loop gain setting procedure, I can only reach about -14 on meter error reading when fine tune is 50 units CW of centre frequency.  My OCXO doesn't have coarse adjust capability, so the fine tune setting can't be easily made '250'.  With the loop closed, the control voltage at the rear bnc is 0V at centre frequency, and goes to +0.8V at 50 units CW fine tune, and to -0.7V at 50 units CCW, so not quite +1 to -1V.  The error meter reading swings from -14 to +14 respectively.  With the fine tune back at 50 units CW, and loop operate switched on, the error falls to 0.

The next adjustment of meter 2nd harmonic level is ok, although the meter needle has some fine jitter of about +/-1 about 35 setting, and A7 TP1 is still about 100mVpp.

Control is now at +10 meter and +0.8Vdc rear bnc, and the 5065A 5Mhz output is now within 1ppb of the GPSDO.  So no 0.01Hz difference that was observed a few days ago, and on review that seems to be a result of not moving the OCXO fine tune to 50 units CW of centre frequency. 

Any thoughts on why the 2nd H level from A7 TP2 is relatively low?  Perhaps I should wait a week or two and confirm if that level changes. 

I don't have access to a frequency standard but will now try and see if the scope phase shift of 5065A to GPSDO can be minimised using the magnetic field adjustment 10 turn pot, given the op manual indicates that adjustment has a nominal 2ppb span (increasing the pot increases the 5065A frequency).

Ciao, Tim

Well I'll be damned !  Now at 400mVpp 273Hz on A7 TP2 - I even had to reset the 2nd H meter level as it was over-ranging. 

A possible catalyst was that I had to turn the unit off to repair the magnetic coil current supply (failed A15 Q6 dual npn) for a few hours.  So with a goodly level of 2nd H I have just been setting the magnetic 'clock' for negligible drift against the GPSDO - down to a level where the GPSDO tracking jitter is circa 0.1ppb (https://www.eevblog.com/forum/projects/budget-gpsdo-a-work-in-progress/50/. 

So I think I am pretty much done with my repair log - touch wood.  Thanks so much to Corby for the tireless assistance, and to Stanley for helping with the heater wire supply.

Ciao, Tim


--- Quote from: trobbins on November 23, 2022, 12:02:52 pm ---A possible catalyst was that I had to turn the unit off to repair the magnetic coil current supply (failed A15 Q6 dual npn) for a few hours.

--- End quote ---

Same failure fixed on my unit:

Best, Stephano

Ta Stephano - I breezed through that thread at the start of the journey, but one then gets blinkered by the immediate faults being faced and forgets to sweep through all the threads again - and yes it makes sense that the correct line wasn't being used given the abnormal scope waveform I had when checking phasing.  In this case the repairs were being done iteratively as functions were being checked, and the magnetic field repair was the last function to check and repair.  And as the ops manual points out, the loss of magnetic field is not monitored by the Continuous operation indicator and logic.  I really should go back and redo all the setup procedure again, but that can be for another day by the owner to enjoy  :)


[0] Message Index

[*] Previous page

There was an error while thanking
Go to full version
Powered by SMFPacks Advanced Attachments Uploader Mod